### v3

# Learning Personalized Decision Support Policies

Abstract
Individual human decision-makers may benefit from different forms of support to improve decision outcomes, but when will each form of support yield better outcomes? In this work, we posit that personalizing access to decision support tools can be an effective mechanism for instantiating the appropriate use of AI assistance. Specifically, we propose the general problem of learning a decision support policy that, for a given input, chooses which form of support to provide to decision-makers for whom we initially have no prior information. We develop Modiste, an interactive tool to learn personalized decision support policies. Modiste leverages stochastic contextual bandit techniques to personalize a decision support policy for each decision-maker. In our computational experiments, we characterize the expertise profiles of decision-makers for whom personalized policies will outperform offline policies, including population-wide baselines. Our experiments include realistic forms of support (e.g., expert consensus and predictions from a large language model) on vision and language tasks. Our human subject experiments add nuance to and bolster our computational experiments, demonstrating the practical utility of personalized policies when real users benefit from accessing support across tasks.
1 Introduction
Human decision-makers use various forms of support to inform their opinions before making a final decision (Keen 1980). Decision-makers with differing expertise may benefit from different forms of support on a given input (Yu et al. 2024). For example, one radiologist may provide a better diagnosis of a chest X-ray by leveraging model predictions (Kahn Jr 1994) while another may perform better after viewing suggestions from senior radiologists (Briggs et al. 2008): see Figure 1. In this paper, we study how to improve decision outcomes by personalizing which form of support we provide to a decision-maker on a case-by-case basis.
Figure 1: Depending on the input, decision-makers need different forms of decision support to make correct decisions. Modiste personalizes access to the right form of support at the right time for the right decision-maker online. Here, Alice would not benefit from model access, while Bob would not benefit from a senior consult.
Since artificial intelligence (AI) is increasingly used as a form of decision support (Lai et al. 2023), even moving towards systems that could act as “thought partners” (Collins et al. 2024b), responsible machine learning (ML) model deployment requires clarity on who should have access to model outputs and when model outputs can be safely exposed to decision-makers (Amodei et al. 2016). Regulation increasingly calls for the “effective and appropriate use” of AI (Biden 2023), requiring careful consideration of when models ought to be accessible to decision-makers. In this paper, we formalize learning a decision support policy that dictates for each individual decision-maker when additional support (e.g., LLM output) should be viewed and used for a given input.
While prior work has assumed access to offline human decisions under support (Laidlaw and Russell 2021; Charusaie et al. 2022) or oracle queries of human behavior (De-Arteaga, Dubrawski, and Chouldechova 2018; Mozannar and Sontag 2020) to learn decision support policies, we argue that this data is unrealistic to obtain in practice across all available forms of support for a new decision-maker. Thus, for individuals for whom we have no prior information initially, we propose learning how to personalize support online. We develop Modiste, [^1] While a “modiste” usually refers to someone who tailors clothing and makes dresses/hats, we use the term to capture our tool’s ability to alter a policy to a decision-maker. an interactive tool that leverages off-the-shelf stochastic contextual bandit algorithms  (Li et al. 2010) to learn decision support policies by modeling human prediction error under varying forms of support.
Our computational experiments explore the utility of personalization across multiple expertise profiles. Based on these experiments, we characterize decision-maker expertise into profiles where personalized policies outperform offline ones, such as population-wide majority vote. We demonstrate that if there is no benefit of personalization, Modiste recovers the same performance as the best form of support.
To validate Modiste on real users ($N=80$), we conduct human subject experiments, where we explore forms of support that include expert consensus, outputs from an LLM, or predictions from a classification model. In contrast to prior work that only tests offline policies or evaluates in simulation, we demonstrate how Modiste can be used to learn personalized decision support policies online on both vision and language tasks. We emphasize our main contributions:
1. Formalizing decision support policies. We propose a formulation for learning a personalized decision support policy that selects the form of support that maximizes a given decision-maker’s performance. We introduce Modiste, a tool to instantiate our formulation using existing methods from stochastic contextual bandits to model human prediction error under different forms of support. We open-source Modiste as a tool to encourage the adoption of personalized decision support policies.
2. Evaluating personalized policies in realistic settings. We use Modiste to learn personalized policies for new decision-makers through both computational and human subject experiments on vision and language tasks. We characterize under which settings we would expect personalized policies to improve performance. Our human subject experiments, where real users interact with Modiste, nuance our findings from the computational experiments on synthetic decision-makers, demonstrating the appropriate use of decision support has benefits in practice.
2 Related Work
Regulating AI Use.
Our study of decision support policies has implications for safely deploying ML models to interact with users. This topic is of heightened importance, particularly in light of recent calls for the “effective and appropriate use” of AI in US President Biden’s Executive Order (Biden 2023) and for clarity on when to “decide not to use [an] AI system” per the EU AI Act (EUA 2023). The disuse of AI can caution downstream misuse of models for assistive decision-making (Brundage et al. 2018). The refusal to use AI assistance can be strategic to empower decision-makers, thus preventing their overreliance on models and encouraging their agency on the task at hand (Gordon and Mugar 2020; Barabas 2022). Each decision-maker may require a different level of use to promote effective use of AI assistance in their decision-making (Kirk et al. 2024); for instance, experts and novices may prefer LLM access in different settings on theorem proving tasks (Collins et al. 2024a). Our experiments engage with such settings by learning when to provide LLM support for language-based tasks via a personalized policy.
Decision Support.
While various forms of decision support have been proposed, such as expert consensus (Scheife et al. 2015) and changes to machine interfaces (Roda 2011), more recent forms of support focus on algorithmic tools where decision-makers are aided by machine learning (ML) models (Phillips-Wren 2012; Gao et al. 2021; Bastani, Bastani, and Sinchaisri 2022). In some prior work, the human does not always make the final decision, such as those that learn to defer decisions from a model to a single decision-maker (Madras, Pitassi, and Zemel 2018; Mozannar and Sontag 2020) or others that jointly learn an allocation function between a model and a pool of decision makers (Keswani, Lease, and Kenthapadi 2021; Hemmer et al. 2022). In our setting, the human is always the decision-maker, which includes settings where humans make the final decisions with support from ML models (Green and Chen 2019; Lai et al. 2023), as well as those where humans make decisions when provided with additional information beyond a model prediction, e.g., explanations (Bansal et al. 2021b), uncertainty (Zhang, Liao, and Bellamy 2020), conformal sets (Babbar, Bhatt, and Weller 2022). While these studies always show a single form of support, recent works consider adapting when AI support is shown to users with a fixed policy. Ma et al. (2023) fit a decision tree to offline user’s decisions to decide when to show AI support to users, and Buçinca et al. (2024) use offline reinforcement learning to estimate if AI support would be helpful, especially under time constraints (Swaroop et al. 2024). Our work considers general forms of support, beyond when to show AI support, and formalizes learning in which contexts each form of support should be provided to an unseen decision-maker online. An extensive comparison to prior work is in the Appendix.
Prior Assumptions About Decision-Maker Information.
We briefly survey the assumptions made about the decision-maker when learning decision support policies. The model of the decision maker is either synthetic, thus lacking grounding in actual human behavioral data, or learned from a batch of offline annotations (Madras, Pitassi, and Zemel 2018; Okati, De, and Rodriguez 2021; Charusaie et al. 2022; Gao et al. 2023). For a new decision-maker or a new form of support, this set of data would not be available in practice. Instead, we propose to learn a decision support policy online to circumvent these limitations. Few works use some aspect of online learning for different decision-making settings or under strict theoretical conditions, as we describe in the Appendix.
Figure 2: We illustrate the process of learning a decision support policy $\pi_{t}$ online to improve a decision-maker $h$’s performance. Since assuming access to sufficient amounts of offline data is unreasonable in practice, our formulation learns a personalized policy online; each decision-maker’s learned policy may differ from that of another decision-maker if they have decisions ($\tilde{y}$) and thus different expertise.
3 Preliminaries
We consider a human decision-making process with different forms of decision support. In our setting, decision-makers may be shown support when selecting an outcome from a fixed set of labels (Seger and Peterson 2013; Lai et al. 2023).
General Problem Formulation.
Decision-makers perform a classification task in observation/feature space ${\mathcal{X}}\subseteq\mathbb{R}^{p}$ and outcome/label space ${\mathcal{Y}}=\left[{K}\right]$. We operate in a stochastic setting where the data $(x,y)\in{\mathcal{X}}\times{\mathcal{Y}}$ are drawn iid from a fixed, unknown data generating distribution ${\mathcal{P}}$, an assumption that reflects typical decision-making settings (Bastani and Bayati 2020; Bastani, Bastani, and Sinchaisri 2022). Importantly, we consider an action set ${\mathcal{A}}$ corresponding to the forms of support available, which may consist of an individual piece of information (e.g., model prediction) or a particular combination of multiple pieces of information (e.g., model prediction and explanation). Given an observation $x\in{\mathcal{X}}$, the human attempts to predict the corresponding label $y\in{\mathcal{Y}}$ using the support prescribed by an action $a\in{\mathcal{A}}$. Note, we do not make assumptions on the specific forms of support $a$, but we provide multiple instantiations in our experiments. The quality of predictions is measured using a 0-1 loss function, where $\ell(y,y^{\prime})=1$ for $y\neq y^{\prime}$ and $\ell(y,y^{\prime})=0$ for $y=y^{\prime}$.
Decision-Making Protocol.
A personalized decision support policy $\pi:{\mathcal{X}}\to\Delta({\mathcal{A}})$ outputs a form of support for a given input. Let $\Pi$ denote the class of all stochastic decision support policies. Let ${\mathcal{A}}=\left\{{\textsc{A}_{1},\ldots,\textsc{A}_{k}}\right\}$, and $\pi(x)_{\textsc{A}_{i}}$ denote $\mathbb{P}[\textsc{A}_{i}\sim\pi(x)]$ for each $\textsc{A}_{i}\in{\mathcal{A}}$. When the policy $\pi$ prescribes the support $\textsc{A}_{i}$, the human decision-maker makes the prediction $\widetilde{y}$ based on the observation $x$ and support $\textsc{A}_{i}$, i.e., the final prediction $\widetilde{y}$ is given by $\widetilde{y}=h(x,\textsc{A}_{i})$. The human decision-making process with different forms of support is described below. For $t=1,2,\dots,T$:
  1. 1.
    A data point $(x_{t},y_{t})\in{\mathcal{X}}\times{\mathcal{Y}}$ is drawn iid from ${\mathcal{P}}$.
  2. 2.
    A form of support $a_{t}\in{\mathcal{A}}$ is selected using a decision support policy $\pi_{t}:{\mathcal{X}}\to\Delta({\mathcal{A}})$.
  3. 3.
    The human decision-maker makes the final prediction $\widetilde{y}_{t}=h(x_{t},a_{t})$ based on $x_{t}$ and $a_{t}$.
  4. 4.
    The human decision-maker incurs a loss $\ell(y_{t},\widetilde{y}_{t})=1$ if $y_{t}\neq\widetilde{y}_{t}$ and $\ell(y_{t},\widetilde{y}_{t})=0$ otherwise.
Evaluation of $\pi$ via Expected Loss.
The quality of a policy $\pi$ can be evaluated using the expected loss incurred by the decision-maker across the input space:
$\begin{split}L_{h}(\pi)&~{}=~{}\mathbb{E}_{(x,y)\sim{\mathcal{P}}}\big{[}\mathbb{E}_{\textsc{A}_{i}\sim\pi(x)}[\ell(y,h(x,\textsc{A}_{i}))]\big{]}.\end{split}$ (1)
We distinguish this metric from the more standard notion of regret, which is typically used to analyze policies in an online learning setting (Li et al. 2010); however, we cannot realize $\pi^{*}$ for an unseen decision-maker in practical scenarios. Thus, we rely on $L_{h}(\cdot)$ as a proxy metric for evaluating the effectiveness of $\pi$.
4 Modiste: Learning Personalized Decision Support Policies
We introduce Modiste, a tool to translate our problem formulation into an interactive interface for learning personalized policies. The workflow, outlined in Figure 2, comprises a learning component to update the personalized policy and an interface to customize the appropriate form of support for each input and each decision-maker.
Learning Problem
To model the decision-making process of a human decision-maker without access to their previous decisions, we consider a stochastic contextual bandit set-up, where the forms of support are the arms, ${\mathcal{X}}$ is the context space, and the policy $\pi$ can be learned online. In Algorithm 1, we detail an algorithm for learning a policy $\pi^{*}$ that minimizes expected loss online. Our goal is to find an optimal decision support policy $\pi^{*}$ that minimizes $L_{h}(\pi)$. We can rewrite Eq. 1 as follows:
$L_{h}(\pi)=\mathbb{E}_{x}\big{[}\sum_{i=1}^{k}{\pi(x)_{\textsc{A}_{i}}\cdot r_{\textsc{A}_{i}}(x;h)}\big{]},$
where $r_{\textsc{A}_{i}}(x;h)=\mathbb{E}_{y|x}[\ell(y,h(x,\textsc{A}_{i}))]$ is the human prediction error for input $x$ and support $\textsc{A}_{i}$. Then, it can be shown that the optimal policy takes the form $\pi^{*}(x)=\operatorname*{arg\,min}_{\textsc{A}_{i}\in\mathcal{A}}r_{\textsc{A}_{i}}(x;h)$: see derivation in Appendix. For Modiste to run, we must maintain and update our estimate of human prediction error $r_{\textsc{A}_{i}}(x;h)$ (Step 1 in Algorithm 1) and of our policy $\pi$ (Step 2 in Algorithm 1).
To update the estimate of human prediction error (Step 1), Modiste implements two approaches to estimate $r_{\textsc{A}_{i}}(x;h)$ for all $x\in{\mathcal{X}}$ and $\textsc{A}_{i}\in{\mathcal{A}}$, but note that any online learning algorithm can be used to update the ${\mathcal{U}}_{r}$. We first consider LinUCB (Li et al. 2010), a common online learning algorithm that approximates the expected loss $r_{\textsc{A}_{i}}(x;h)$ by a linear function $\widehat{r}_{\textsc{A}_{i}}(x;h):=\left\langle{\theta_{\textsc{A}_{i}}},{x}\right\rangle$. Although the linearity assumption may not hold in general, we learn the parameters $\left\{{\theta_{\textsc{A}_{i}}:\textsc{A}_{i}\in{\mathcal{A}}}\right\}$ using LinUCB with the instantaneous reward function $R(x,y,\textsc{A}_{i};h):=-\ell(y,h(x,\textsc{A}_{i}))$. We then normalize the resulting $\widehat{r}_{\textsc{A}_{i}}(x;h)$ values to lie in the range $[0,1]$. The second algorithm we use is an intuitive $K$-nearest neighbor (KNN) approach, which is a simplified variant of KNN-UCB (Guan and Jiang 2018). Here, we maintain an evolving data buffer ${\mathcal{D}}_{t}$, which accumulates a history of interactions with the decision-maker. For any new observation $x$, we estimate $\widehat{r}_{\textsc{A}_{i}}(x;h)$ values by finding $K$-nearest neighbors in ${\mathcal{D}}_{t}$ and computing the average error of these neighbors.
Algorithm 1 Learning a decision support policy
1:  Input: human decision-maker $h$
2:  Initialization: data buffer ${\mathcal{D}}_{0}=\{\}$; human error values $\{\widehat{r}_{\textsc{A}_{i},0}(x;h)=0.5:x\in{\mathcal{X}},\textsc{A}_{i}\in{\mathcal{A}}\}$; initial policy $\pi_{1}$
3:  for $t=1,2,\dots,T$ do
4:     data point $(x_{t},y_{t})\in{\mathcal{X}}\times{\mathcal{Y}}$ is drawn iid from ${\mathcal{P}}$
5:     support $a_{t}\in{\mathcal{A}}$ is selected using policy $\pi_{t}$
6:     human makes the prediction $\widetilde{y}_{t}$ based on $x_{t}$ and $a_{t}$
7:     human incurs the loss $\ell(y_{t},\widetilde{y}_{t})$
8:     update the buffer ${\mathcal{D}}_{t}\leftarrow{\mathcal{D}}_{t-1}\cup\left\{{(x_{t},a_{t},\ell(y_{t},\widetilde{y}_{t}))}\right\}$
9:     update the decision support policy:
$\widehat{r}_{\textsc{A}_{i},t}(x;h)~{}\leftarrow~{}$ ${\mathcal{U}}_{r}(\widehat{r}_{\textsc{A}_{i},t-1}(x;h),{\mathcal{D}}_{t}),\quad\forall\textsc{A}_{i}\in\mathcal{A}$ (Step 1)
$\pi_{t+1}(x)~{}\leftarrow~{}$ ${\mathcal{U}}_{\pi}(\left\{{\widehat{r}_{\textsc{A}_{i},t}}\right\}_{i})$ (Step 2)
10:  end for
11:  Output: policy $\pi_{\lambda}^{\text{alg}}\leftarrow\pi_{T+1}$
In practical settings where interactions are limited (like in our human subject experiments), the number of interactions $T$ tends to be relatively small, which renders pure exploratory policies infeasible (Sutton and Barto 2018). Thus in Step 2, we guide exploration of the policy via:
$\pi_{t+1}(x)=\operatorname*{arg\,min}_{\textsc{A}_{i}\in\mathcal{A}}\widehat{r}_{\textsc{A}_{i},t}(x;h)+b_{\textsc{A}_{i},t}(x;h),$
where $b_{\textsc{A}_{i},t}(x;h)$ corresponds to some exploration bonus. In the Appendix, we provide implementations of Modiste with LinUCB and with KNN.
Modiste Interface
We provide an extendable interface for the study and deployment of decision-support policies. At each time step, Modiste sends each user’s predictions to a server running Algorithm 1, which identifies the next form of support for the next input. Modiste then updates the interface accordingly to reflect the selected form of support. Our tool can be flexiblely linked to crowdsourcing platforms like Prolific (Palan and Schitter 2018). We implement three common forms of support: (1) Human Alone, where the human makes the decision solely based on the input, (2) Model Prediction, which shows decision-makers a model’s prediction for the given input (Bastani, Bastani, and Sinchaisri 2022), and (3) Expert Consensus, which presents the user with a distribution over labels from multiple annotators (Scheife et al. 2015). In Figure 3, we provide an example screenshot of the interface. Participants are informed of their own correctness after each trial and the correctness of the form of support (e.g., model prediction) if support was provided, so that participants can learn whether support ought to be relied upon.
Figure 3: Example of the Modiste interface for MMLU-$2A$ where the human is provided responses from a LLM.
5 Experimental Set-up
Before we evaluate Modiste, we overview the set-up of subsequent experiments, both computational and human subject. All other details are in the Appendix.
Decision-making Tasks
Following prior studies on human-AI interactions (Babbar, Bhatt, and Weller 2022; Mozannar et al. 2023; Lee et al. 2023), the decision-making tasks in our experiments center around the following vision and language datasets:
  1. 1.
    CIFAR-10 (Krizhevsky 2009), a 10-class image classification dataset;
  2. 2.
    MMLU (Hendrycks et al. 2020), a multi-task text-based benchmark that tests for knowledge and problem-solving ability across $57$ topics in both the humanities and STEM.
In terms of the size of $|\mathcal{A}|$, we let $kA$ denote when there are $k$ forms of support for a task. We focus on $k=2$ or $3$, which captures a buffet of real-world scenarios in prior work where decision-makers have a few tools at their disposal. In particular, the two action setting covers practical use cases where a decision-maker has the option of using a model or not. The learned decision support policy then reflects appropriate use, as the model would be hidden when a decision-maker does not need it. While the forms of support we consider are common in practice (Lai et al. 2023), our choices of support are not intended to exhaustively demonstrate the diverse forms of support that Modiste can handle. We now describe our two main tasks, which are designed to be accessible to crowdworkers and will be featured in both the computational and human subject experiments. [^2] In the Appendix, we include computational experiments for two additional tasks (Synthetic-$2A$ and CIFAR-$2A$), and experiments where we vary the size of $k$.
CIFAR-$3A$.
In this task, we consider three forms of support: Human Alone, Model, Consensus. Our goal is to construct a setup reflecting a realistic setting in which different forms of support result in different strengths and weaknesses for decision-makers. To instantiate this setting, we deliberately corrupt images of different classes to evoke performance differences – necessitating that a decision-maker appropriately calibrate when to rely on each form of support. We consider $5$ of the animal classes in CIFAR-10; of these, we never corrupt images of Birds, do not corrupt images of Deers and Cats for the Model, and do not corrupt images of Horses and Frogs for the Consensus.
MMLU-$2A$.
The two forms of support are Human Alone and LLM, where the human is provided responses generated from InstructGPT3.5, text-davinci-003(Ouyang et al. 2022) using the same few-shot prompting scheme for MMLU as Hendrycks et al. (2020). We conducted pilot studies to select a subset of topics where the accuracy of the LLM and average human accuracy vary. We choose the following topics: Computer Science, US Foreign Policy, High School Biology, and Elementary Mathematics. The goal of this task is to evaluate whether we can learn personalized support “in-the-wild,” where we naturally expect people to excel at different topics, akin to real-world settings where decision-makers may have varying expertise.
Baselines and Other Parameters
Algorithms and Baselines. We compare personalized policies, learned using Algorithm 1 with LinUCB and with KNN reporting results as Modiste-LinUCB and Modiste-KNN respectively, against the following offline policies:
  • Human + Support, where the decision-maker always receives the same form of support: $\pi(x)=A_{i}$ for all $x$. In CIFAR-$3A$, there are 3 fixed support baselines, corresponding to each form of support. In MMLU-$2A$, there are 2 fixed support baselines.
  • Population-wide, where the decision-maker receives a form of support based on the majority vote from 10 learned policies (breaking ties at random). For this baseline, the form of support may vary across contexts but is not personalized to individual needs. This baseline is akin to recent offline policy learning (Ma et al. 2023; Buçinca et al. 2024).
Number of Interactions. While more interactions (higher $T$) provide more data points to estimate each $r_{\textsc{A}_{i}}$, we need to consider what a realistic value of $T$ is given constraints of working with real humans (e.g., limited attention and cognitive load). In online learning, $T$ is usually unreasonably large, on the order of thousands (Li et al. 2010; Guan and Jiang 2018). Via pilot studies, we found that 100 CIFAR images or 60 MMLU questions were a reasonable number of decisions to make within 20-40 minutes (a typical time limit for an online study), which we use throughout our experiments.
Algorithm Invariant Strictly Better Varying
H-Only $0.00\pm 0.01$ $0.09\pm 0.08$ $0.50\pm 0.06$
H-Model $0.00\pm 0.01$ $0.22\pm 0.19$ $0.35\pm 0.05$
H-Consensus $0.00\pm 0.01$ $0.23\pm 0.13$ $0.27\pm 0.08$
Population $0.00\pm 0.02$ $0.18\pm 0.08$ $0.15\pm 0.03$
Modiste-LinUCB $0.00\pm 0.01$ $0.17\pm 0.05$ $0.19\pm 0.05$
Modiste-KNN $0.00\pm 0.01$ $0.06\pm 0.01$ $\mathbf{0.08\pm 0.02}$
Algorithm Invariant Strictly Better Varying
H-Only $0.01\pm 0.01$ $0.18\pm 0.17$ $0.22\pm 0.12$
H-LLM $0.01\pm 0.01$ $0.18\pm 0.21$ $0.12\pm 0.17$
Population $0.00\pm 0.02$ $0.19\pm 0.07$ $0.12\pm 0.09$
Modiste-LinUCB $0.00\pm 0.01$ $0.12\pm 0.03$ $0.07\pm 0.04$
Modiste-KNN $0.01\pm 0.01$ $0.05\pm 0.03$ $\mathbf{0.05\pm 0.03}$
Table 1: We evaluate Modiste across three expertise profiles. We compute the average excess loss $L_{h}(\pi)-L_{h}^{opt}$ (lower is better), and standard deviation across individuals in each expertise profile for both CIFAR-$3A$ (Left) and MMLU-$2A$ (Right). $L_{h}(\pi)$ is computed by averaging across the last 10 steps of 100 total time steps. We bold the variant with the lowest excess loss that is statistically significant from the other variants. Note that this only occurs in the “varying” expertise setting.
6 Computational Evaluation
Decision-makers may have different “expertise” (i.e., strengths and weaknesses) across input space $\mathcal{X}$ under each form of support. We evaluate Modiste using simulated human behavior to capture diverse decision-maker expertise.
Expertise Profiles
We can capture an individual $h$’s expertise via an expertise profile, which is defined over the input space ${\mathcal{X}}$. We divide ${\mathcal{X}}$ into disjoint regions (i.e., ${\mathcal{X}}=\cup_{j\in[N]}{\mathcal{X}}_{j}$); these regions could be defined by class labels or by covariates, depending on the task. [^3] While we instantiate decision-makers this way, Modiste does not take the expertise profiles or how they were constructed (e.g., the regions) as input. We let $r_{\textsc{A}_{i}}({\mathcal{X}}_{j};h)$ denote $h$’s average prediction error under support $\textsc{A}_{i}$ across region ${\mathcal{X}}_{j}$.
Human-informed synthetic decision-makers.
To construct expertise profiles with realistic values for each $r_{\textsc{A}_{i}}({\mathcal{X}}_{j};h)$, we collect data on user decisions across different users and then calculate individual $r_{\textsc{A}_{i}}({\mathcal{X}}_{j};h)$. The set of participant expertise profiles form a population of decision-makers that we refer to as human-informed synthetic decision-makers. From the estimated $r_{\textsc{A}_{i}}({\mathcal{X}}_{j};h)$ of each human-informed synthetic decision-maker, we can simulate decision-maker behavior.
To construct human-informed synthetic decision-makers, we recruited 20 participants from Prolific (10 for CIFAR-$3A$ and 10 for MMLU-$2A$). We use the same recruitment scheme as the larger human subject experiment described in the Appendix. We define regions of expertise over class labels for CIFAR-$3A$ and over question topics for MMLU-$2A$, as we expect $r_{\textsc{A}_{i}}(x;h)$ to be roughly constant for $x\in{\mathcal{X}}_{j}$ where ${\mathcal{X}}_{j}$ is defined by a class label or question topic. We showed each participant similar inputs with different forms of support to estimate $r_{\textsc{A}_{i}}({\mathcal{X}}_{j};h)$ for each support $\textsc{A}_{i}$ in each region ${\mathcal{X}}_{j}$. On each trial, each participant is randomly assigned a form of support; trials are approximately balanced by the type of support and grouping (i.e., topic or class). We compute participant accuracy averaged over all trials: 100 for CIFAR-$3A$, 60 for MMLU-$2A$. We denote expertise profiles as follows: if there were three regions in the input space, an individual’s expertise profile under support $\textsc{A}_{i}$ would be written as $r_{\textsc{A}_{i}}=[0.7,0.1,0.7]$, meaning the individual incurs a loss of $0.7$ on ${\mathcal{X}}_{1}$, $0.1$ on ${\mathcal{X}}_{2}$, and $0.7$ on ${\mathcal{X}}_{3}$.
Policies for each profile.
Based our pilot study, we define the three expertise profiles and what kind of decision support policy we expect to be learned for each:
  • Approximately Invariant expertise across all the regions under different forms of support, i.e., $r_{\textsc{A}_{1}}({\mathcal{X}}_{j};h)\approx r_{\textsc{A}_{2}}({\mathcal{X}}_{j};h)\approx\cdots\approx r_{\textsc{A}_{k}}({\mathcal{X}}_{j};h),\forall j\in[N]$. Both a random decision support policy and a policy that selects a fixed form of support would suffice for such a profile.
  • Varying expertise where a decision-maker excels in some areas but benefits from support in areas beyond their training (Schvaneveldt et al. 1985), i.e., $r_{\textsc{A}_{1}}({\mathcal{X}}_{j};h)\leq r_{\textsc{A}_{2}}({\mathcal{X}}_{j};h)\text{ and }r_{\textsc{A}_{2}}({\mathcal{X}}_{k};h)\leq r_{\textsc{A}_{1}}({\mathcal{X}}_{k};h),\text{for some }j,k\in[N]$. For this expertise profile, we expect the decision support policy to select different forms of support in different regions. The quantity of $|r_{\textsc{A}_{1}}$ $r_{\textsc{A}_{2}}|$ for a region will dictate how efficiently the policy can be learned.
  • Strictly Better expertise (e.g., $\textsc{A}_{1}\succ\textsc{A}_{2}\succ\cdots\succ\textsc{A}_{k}$) that is uniformly maintained across all the regions, i.e., $r_{\textsc{A}_{1}}({\mathcal{X}}_{j};h)\leq r_{\textsc{A}_{2}}({\mathcal{X}}_{j};h)\leq\cdots\leq r_{\textsc{A}_{k}}({\mathcal{X}}_{j};h),\forall j\in[N]$. A decision support policy should learn the fixed form of support to use for all inputs.
Per the task design, we find participants generally only display varying expertise profiles on CIFAR-$3A$ while we find instances of all three expertise profiles on MMLU-$2A$.
When is Personalization Useful?
We investigate how personalized policies compare against offline baselines under each expertise profile (Table 1). We verify that learning decision support policies are not helpful for decision-makers with “invariant” expertise profiles. For individuals who fall under the “varying” profiles, we find at least one personalized policy outperforms offline policies and learns a policy that is significantly closer to the decision-maker’s optimal performance. This is because a personalized policy identifies which form of support is better in each context, compared to fixed offline policies which always show one form of support or to the population-wide variant, which may not provide the correct form of support to each individual. For individuals in the “strictly better” profile, while we do not find a statistically significant difference from the baselines due to the large variance in fixed policies (e.g., they work well for some decision-makers but poorly for others), we observe that much smaller variance with Modiste, particularly using KNN. For most individuals, the population-wide baseline performs poorly, emphasizing the need for personalization of decision support. Misalignment between the population policy and the optimal policy of the new decision-maker demonstrably leads to an ineffective use of decision support. We note that KNN generally outperforms LinUCB, the latter of which can be saddled by its implicit linearity assumption. We further study the effect of various parameters, e.g., exploration parameters, KNN parameters, embedding size, and the number of interactions in the Appendix.
7 Modiste with Real Users
To validate whether Modiste can improve decision-maker performance in practice, we run a series of human subject experiments (i.e., ethics-reviewed studies with real human participants). We first introduce the set-up of our user study; additional information can be found in the Appendix.
Recruitment.
We recruit a total of $80$ crowdsourced participants from Prolific to interact with Modiste ($N=30$ and $N=50$ for CIFAR-$3A$ and MMLU-$2A$, respectively). We recruit more participants for MMLU-$2A$, as we expect greater individual differences in regions where support is needed, e.g., some participants may be good at mathematics and struggle in biology, whereas others may excel in biology.
Each participant is assigned to only one task. Within a task, participants are randomly assigned to one algorithm variant; an equal number of participants are included per variant (i.e., 10 for MMLU and 5 for CIFAR). Participants are required to reside in the United States and speak English as a first language. Participants are paid at a base rate of \$9/hr, and are told they may be paid an optional bonus up to \$10/hr based on the number of correct responses. We allot 25-30 minutes for the CIFAR task and 30-40 for MMLU, as each MMLU question takes more effort. We applied the bonus to all participants in all studies. We run an ANOVA with Tukey HSD across the conditions for each task.
Modiste outperforms baselines for “varying” expertise profiles.
By design, the CIFAR-$3A$ task compels “varying” profiles: Modiste’s forte. This is reflected quantitatively in Figure 4 (left), where both Modiste variants have lower expected losses than any of the offline policies. In particular, KNN achieves statistically significantly lower expected loss over all variants ($p<0.001$ across all pairs), including the population-wide policy learned from the pilot study. When visualizing the learned decision support policies, we observe that Modiste can reconstruct near-optimal policies, as depicted in the Appendix.
Figure 4: We report expected average loss $L_{h}(\pi)$ (lower is better) and standard error in the last 10 trials by Prolific participants for each algorithm, with CIFAR conditions on the left and MMLU conditions on the right. In the CIFAR setting, where individuals typically exhibit “varying” expertise profiles, we see significant benefits from using Modiste, particularly in the KNN setting. While we observe that most individuals in the MMLU condition exhibit “strictly better” expertise, which means personalized policies typically only perform as well as the best baseline, we still observe instances of deferred decisions to the human on a case-by-case basis—see Figure 5.
Figure 5: Snapshots of the learned decision support policies computed at the end of the study for 10 participants on the MMLU task. The forms of support are colored in t-SNE embedding space. All participants exhibit distinct policies across input space. The bar plot to the right of each scatter plot shows the relative performance of that decision-maker alone in each category, ordered from left to right as M=Mathematics, B=Biology, CS=Computer Science, FP=Foreign Policy per subplot. When a decision-maker performs well alone, Modiste learns policies to empower that decision-maker without LLM access. For example, the individual in the top left is highly competent at both Mathematics and Foreign Policy; the learned decision support policy reflects this.
Modiste matches the best baseline for “strictly better” profiles.
Polymaths are rare; we observe no different in our human subject experiments on MMLU, as most participants are better with LLM access, which places them in the “strictly better” category. While participants with both Modiste variants outperform the H-Only baseline ($p<0.01$), we observe that on average Modiste settings are no different than H-LLM, the condition where people always have access to the LLM, as shown in Figure 4 (right). This confirms our hypothesis from the computational experiments. Further, while the average performance of Modiste variants is similar to that of the offline fixed or population-wide policies in Figure 4 (right), the variance is significantly smaller—particularly with KNN.
Modiste can facilitate appropriate use of decision support.
Since the LLM only excels at three of the four topics on the MMLU task, Modiste learns in many cases to defer to human judgment on Mathematics questions, particularly when the human has strong expertise. In Figure 5, we visualize the learned decision support policies of various individuals in the study and illustrate how Modiste yields policies that provide support on different topics for different decision-makers.
Limitations.
In this work, we consider the classification setting where we get immediate feedback (e.g., we can calculate the loss to update $\pi$). Future work can consider more complex decision-making tasks that may require extending to a delayed feedback setting or to a different cognitive task (e.g., planning or perception). Though Modiste is promising, we note that significant issues can arise when decision-makers blindly rely on decision support (Buçinca et al. 2020; Chen et al. 2023), especially when the support is erroneous or ineffective; such over-reliance requires careful attention to prevent. Further, our problem definition hinges on domain experts defining the available forms of support (i.e., we need a clearly defined $\mathcal{A}$ to use Modiste). In practice, this may prove difficult, as one may not know how to define specific forms of support or decision-makers may have access to varying support sets.
8 Conclusion
A decision support policy captures when and which form of support should be provided to improve a decision-maker’s performance. The selective use of AI-based decision support helps instantiate the “appropriate use” clauses in emerging regulation (Biden 2023), as we only provide AI assistance to decision-makers as and when it is beneficial to them. We introduce Modiste, an interactive tool for learning a decision support policy for each decision-maker using contextual bandits. To the best of our knowledge, we are the first to learn and validate such a policy online for unseen decision-makers. Within our Modiste interface, we instantiate two variants of Algorithm 1 using existing stochastic contextual bandit tools, namely LinUCB and online KNN. Our computational and human subject experiments highlight the importance—and feasibility—of personalizing decision support policies for individual decision-makers. Our human subject experiments show promise, as we personalize decision support policies in few iterations yet find nuances in decision-makers’ need for support: some unskilled decision-makers uniformly benefit from LLM access, while others only need LLMs for some tasks. While encouraging rich cross-talk between domain experts and practitioners, future work integrating Modiste into existing decision-making workflows would pave a route towards responsible use of AI as decision support.
Acknowledgements
UB acknowledges support from DeepMind and the Leverhulme Trust via the Leverhulme Centre for the Future of Intelligence (CFI). AW acknowledges support from a Turing AI Fellowship under grant EP/V025279/1, EPSRC grants EP/V056522/1 and EP/V056883/1, and the Leverhulme Trust via CFI. KMC acknowledges support from a Marshall Scholarship and the Cambridge Trust. PK acknowledges support from The Alan Turing Institute. AT acknowledges support from the National Science Foundation grants IIS1705121, IIS1838017, IIS2046613, IIS2112471, an Amazon Web Services Award, a Facebook Faculty Research Award, funding from Booz Allen Hamilton Inc., and a Block Center Grant. This work is supported (in part) by ELSA - European Lighthouse on Secure and Safe AI funded by the European Union under grant agreement No. 101070617. Views and opinions expressed are however those of the author(s) only and do not necessarily reflect those of any of these funding agencies, European Union or European Commission. We thank the Prolific participants and our volunteer pilot participants. We thank Alan Liu and Kartik Chandra for their invaluable input on designing our web-based server.
Appendix A Comparison Against Prior Work
Most papers on human-AI collaboration have considered clever ways of abstaining from prediction on specific inputs (Cortes, DeSalvo, and Mohri 2016; Cortes et al. 2018), learning deferral functions based on multiple experts (Vovk 1998; Keswani, Lease, and Kenthapadi 2021), or teaching decision-makers when to rely (Mozannar, Satyanarayan, and Sontag 2022). There are also a number of papers from the HCI literature (see survey by (Lai et al. 2023)) that evaluate the two-action setting of our formulation using a static policy (e.g., always showing the ML model prediction or always showing some form of explanation).
To clarify how our set-up and assumptions differ from prior work, we overview work that we believe could be considered most similar to ours. We decompose our comparisons along a few dimensions: Decision-support set-up: Does the human make the final decision, or is it a different set-up? Assumptions about decision-maker information: What does prior work assume about access to a decision-maker when learning a policy? Evaluation: Does prior work simulate humans? Does prior work run user studies?
Mozannar and Sontag (2020) and Hemmer et al. (2023):
  • Decision-support set-up: This work’s set-up can be considered a two-action setting of our formulation, where $\mathcal{A}=\{\texttt{DEFER},\texttt{MODEL}\}$. Extending the work of Madras, Pitassi, and Zemel (2018), this work proposes the learning to defer paradigm, where the decision-maker may not always make a final decision (i.e., sometimes the decision is deferred entirely to an algorithmic-based system). In our set-up, deferring to a MODEL is equivalent to always adhering to a label-based form of support. The human is always the final decision-maker in our work, which is representative of many decision-making set-ups in practice (Lai et al. 2022), but not captured in this line of prior work.
  • Assumptions about decision-maker information: This work assumes oracle query access to the decision-maker, for whom they are learning a policy.
  • Evaluation: This work evaluates their approach using human simulations (no real human user studies).  Mozannar and Sontag (2020) define synthetic experts in the following way: “if the image belongs to the first $k$ classes the expert predicts perfectly, otherwise the expert predicts uniformly over all classes.”
Gao et al. (2021) and Gao et al. (2023):
  • Decision-support set-up: This work defines two actions: $\mathcal{A}=\{\texttt{DEFER},\texttt{MODEL}\}$. They do not consider the, more practical assumption that the decision-maker will view a model prediction before making a decision themselves. Their formulation is similar to the above but they use offline bandits to learn a suitable policy.
  • Assumptions about decision-maker information: Gao et al. (2021) assume that understanding decision-maker’s expertise (at a population-level, not at a individual-level) can help learn better routing functions (i.e., defer only when appropriate). Gao et al. (2023) assume access to a decision history for each decision-maker.
  • Evaluation: They run a human subject experiment to collect offline annotations, which can be used to learn when to defer to decision-makers. Gao et al. (2023) goes further to personalize a deferral policy based on offline annotations for each decision-maker.
Bordt and Von Luxburg (2022):
  • Decision-support set-up: This work’s set-up can be considered a two-action setting of our formulation, where $\mathcal{A}=\{\texttt{DEFER},\texttt{SHOW}\}$; however, they are concerned with the learnability of such a set up. They do not devise algorithms for this setting, as they are only focused on its theoretical formulation.
  • Assumptions about decision-maker information: They assume the decision-maker has access to information not contained in the input but still important to the task.
  • Evaluation: This is a theory paper, containing neither computational nor human subject experiments.
Noti and Chen (2023):
  • Decision-support set-up: This work’s set-up can be considered as a two-action setting of our formulation, where $\mathcal{A}=\{\texttt{DEFER},\texttt{SHOW}\}$. This is not an online algorithm and as such, the policy does not update.
  • Assumptions about decision-maker information: They assume access to a dataset of human decisions and that all decision-makers are similar (i.e., they deploy one policy for all decision-makers).
  • Evaluation: This work is one of few that runs a user study to evaluate their (fixed) policy on unseen decision-makers.
Babbar, Bhatt, and Weller (2022):
  • Decision-support set-up: This work considers the two-action setting of our formulation, where $\mathcal{A}=\{\texttt{DEFER},\texttt{CONFORMAL}\}$. Their policy is learned offline, is the same for all decision-makers, and is not updated in real-time based on decision-maker behavior.
  • Assumptions about decision-maker information: They use CIFAR-10H (Peterson et al. 2019) to learn a population-level deferral policy. This assumes that we have annotations for each decision-maker for every datapoint and assumes that all new decision-makers have the same expertise profiles as the population average.
  • Evaluation: This work runs a user study to evaluate their (fixed) policy. They show the benefits of DEFER+CONFORMAL over CONFORMAL or SHOW alone.
Wolczynski, Saar-Tsechansky, and Wang (2022):
  • Decision-support set-up: This work considers two actions per our formulation, where $\mathcal{A}=\{\texttt{DEFER},\texttt{SHOW}\}$. They learn a rule-based policy offline for each decision-maker.
  • Assumptions about decision-maker information: They simulate human behavior by considering explicit functions of how human expertise may vary in input space.
  • Evaluation: While this work does consider the human to be the final decision-maker, they only validate their proposal in simulation, not on actual human subjects.
Ma et al. (2023):
  • Decision-support set-up: This work considers two actions per our formulation, where $\mathcal{A}=\{\texttt{DEFER},\texttt{SHOW}\}$. They learn a tree-based and rule-based policy offline for the population but permit one-off editing of these policies.
  • Assumptions about decision-maker information: They assume access to offline data to initialize their policies, which are then editable. They do not automatically learn online from user behavior over future interactions.
  • Evaluation: They run extensive human subject experiments to show the benefit of fixed policies, learned from a population and edited by each decision-maker. On models trained on simple tabular UCI datasets, they qualitatively study a decision-maker’s trust and perception of assistance.
Buçinca et al. (2024):
  • Decision-support set-up: This work considers four actions per our formulation, where $\mathcal{A}=\{\texttt{DEFER},\texttt{EXPLAIN},\texttt{CLICK},\texttt{SHOW},\}$. They learn a reinforcement learning policy offline for each decision-maker. This requires having data for each action.
  • Assumptions about decision-maker information: They require a purely exploratory phase in their learning procedure: this prior data can then be used to learn the policy.
  • Evaluation: They run human subject experiments on a toy task for assigning exercises to characters, picking between two classes. This differs from our realistic MMLU setting where optionality is higher as well.
We now list various forms of support that can be included in the action space of our problem formulation. The design of the action space is up to domain experts, who can decide not only which actions are feasible but also how much cost to assign to each form of support.
  • DEFER: This form of support is equivalent to no support. Decision-makers are asked to make a decision without any assistance. The machine learning community has studied how to identify when to defer to a subset of examples to humans based on human strengths (Bansal et al. 2021a; Wilder, Horvitz, and Kamar 2021) and/or model failures (Chow 1957; Geifman and El-Yaniv 2017). The premise of such an action would be to allow human decision-makers to be unaided and squarely placing decision liability on the individual.
  • SHOW: In many settings, machine learning (ML) models are trained to do prediction tasks similar to the decision-making task prescribed to the human, or in the case of foundation models (Bommasani et al. 2021), ML systems can be adapted to aid decision-making, even if the task was not specifically prescribed at train-time (Yang et al. 2023). In essence, a machine learning model prediction, or associated generation (e.g., a code snippet (Mozannar et al. 2022)) would be shown to aid an individual decision-maker. This has been shown to help improve decision-maker performance. The following are variations of showing a model prediction to a decision-maker.
    • CONFORMAL: For classification tasks, only displaying the most likely label may not lead to good performance due to various reasons, including uncertainty in the modeling procedure (Vovk, Gammerman, and Shafer 2005; Bondi et al. 2022); however, such uncertainty can be communicated to decision-maker by showing a prediction set to experts (Babbar, Bhatt, and Weller 2022). Such a prediction set might be generated using conformal prediction, which guarantees the true label lies in the set with a user-specified error tolerance (Bates et al. 2021; Straitouri et al. 2022).
    • CONFIDENCE: Instead of translating the uncertainty into a prediction set (or interval), one could simply show the confidence or uncertainty of the prediction, which may manifest as displaying probabilities, standard errors, or entropies (Spiegelhalter 2017; Bhatt et al. 2021). The visualization mechanism used for displaying confidence may alter the decision-maker’s performance (Hullman et al. 2018; Zhang, Liao, and Bellamy 2020).
    • EXPLAIN: In addition to providing a model prediction, many have considered showing an explanation of model behavior, examples of which include feature attribution (Ribeiro, Singh, and Guestrin 2016; Buçinca et al. 2020), sample importance (Kim, Rudin, and Shah 2014; Jeyakumar et al. 2020), counterfactual explanations (Ustun, Spangher, and Liu 2019; Antoran et al. 2020), and natural language rationales (Ehsan et al. 2018; Camburu et al. 2018). Displaying such explanations to end users has had mixed results on how decision-making performance is affected (Chen et al. 2022; Lai et al. 2022). Worryingly, in many settings, showing some types of explanations may lead to to over-reliance on models by giving the perception of competence (Buçinca et al. 2020; Zerilli, Bhatt, and Weller 2022; Chen et al. 2023).
  • CONSENSUS: One can also depict forms of support that are independent of any model, for instance, presenting the belief of one or more humans. Belief distributions can be constructed by pooling over many different humans’ “votes” for what a label ought to be, e.g., (Peterson et al. 2019; Beyer et al. 2020; Uma, Almanea, and Poesio 2022; Uma et al. 2020; Gordon et al. 2021, 2022), or by eliciting distributions over the likely label directly from each individual human (Collins, Bhatt, and Weller 2022; Collins et al. 2023). These consensus distributions permit the expression of uncertainty without any model. However, the elicitation of this form of support may be costly and humans may be fallible in the information they provide, e.g., due to direct labeling errors (Dawid and Skene 1979; Augustin et al. 2017; Whitehill et al. 2009; Wei et al. 2022), or miscalibrated confidence (O’Hagan et al. 2006; Collins et al. 2023; Lichtenstein, Fischhoff, and Phillips 1977; Tversky and Kahneman 1996).
  • ADDITIONAL: While much of this paper focused on support that provides decision-makers with label information, decision support may also entail acquiring or displaying additional contextual information (e.g., new features (Bakker et al. 2021)) or requesting previously unseen features, for instance, through additional medical diagnostics  (Harrell et al. 1982; Mylonakis et al. 2000). This flavor of support can be varied structurally, ranging from the results of a search query (Nakano et al. 2021) to hierarchical information like exposing the subsidiary ownership structure for multinational corporations (Erramilli 1996). Some pieces of additional information may require additional cost, or certification if pertaining to sensitive attributes.
Prior work on multi-objective contextual bandits.
We summarize why some theoretical work on multi-objective contextual bandits cannot be directly applied to our problem formulation. In their study,  Tekin and Turğay (2018) addressed a contextual multi-armed bandit problem with two objectives, where one objective dominates the other. Their aim was to maximize the total reward in the non-dominant objective while ensuring that the dominant objective’s total reward is also maximized. However, our specific case requires the minimization of the total expected loss of the support policy while ensuring that the expected accuracy of the decision-maker under the policy remains above a certain threshold. Therefore, the techniques presented by here cannot be directly applied to our scenario. Turgay, Oner, and Tekin (2018) investigated the multi-objective contextual bandit problem with similarity information. Their approach relies on the assumption that a Lipschitz condition holds for the set of feasible context-arm pairs concerning the expected rewards for all objectives and that the learning algorithm has knowledge of the corresponding distance function (see Assumption 1 of their paper).
Appendix B Additional Details on Problem Formulation
The optimization problem in the standard setting, where the only objective relates to expected loss, can be written as:
$\min_{\pi\in\Pi}L_{h}(\pi)~{}=~{}$ $\min_{\pi\in\Pi}\mathbb{E}_{(x,y)\sim{\mathcal{P}}}\big{[}\mathbb{E}_{\textsc{A}_{i}\sim\pi(x)}[\ell(y,h(x,\textsc{A}_{i}))]\big{]}$
=$~{}\stackrel{{\scriptstyle(a)}}{{=}}~{}$ $\min_{\pi\in\Pi}\mathbb{E}_{(x,y)\sim{\mathcal{P}}}\bigg{[}\sum_{i=1}^{k}{\pi(x)_{\textsc{A}_{i}}\cdot\ell(y,h(x,\textsc{A}_{i}))}\bigg{]}$
=$~{}\stackrel{{\scriptstyle(b)}}{{=}}~{}$ $\min_{\pi\in\Pi}\mathbb{E}_{x}\bigg{[}\sum_{i=1}^{k}{\pi(x)_{\textsc{A}_{i}}\cdot\mathbb{E}_{y|x}[\ell(y,h(x,\textsc{A}_{i}))]}\bigg{]}$
=$~{}\stackrel{{\scriptstyle(c)}}{{=}}~{}$ $\min_{\pi\in\Pi}\mathbb{E}_{x}\bigg{[}\sum_{i=1}^{k}{\pi(x)_{\textsc{A}_{i}}\cdot r_{\textsc{A}_{i}}(x;h)}\bigg{]},$
where $(a)$ is due to the notation $\pi(x)_{\textsc{A}_{i}}:=\mathbb{P}[\textsc{A}_{i}\sim\pi(x)]$, $(b)$ is due to the properties of expectation, and $(c)$ is due to the notation $r_{\textsc{A}_{i}}(x;h):=\mathbb{E}_{y|x}[\ell(y,h(x,\textsc{A}_{i}))]$. Then, by noting the fact that the expression $\mathbb{E}_{x}\bigg{[}\sum_{i=1}^{k}{\pi(x)_{\textsc{A}_{i}}\cdot r_{\textsc{A}_{i}}(x;h)}\bigg{]}$ can be optimized independently for each $x\in\mathcal{X}$, we can rewrite the above optimization problem as follows, for each $x\in\mathcal{X}$:
$\min_{\pi(x)\in\Delta(\mathcal{A})}\sum_{i=1}^{k}{\pi(x)_{\textsc{A}_{i}}\cdot r_{\textsc{A}_{i}}(x;h)}~{}=~{}\min_{\textsc{A}_{i}\in\mathcal{A}}r_{\textsc{A}_{i}}(x;h).$
Thus, an optimal policy for the above optimization problem is: $\pi^{*}(x)\in\operatorname*{arg\,min}_{\textsc{A}_{i}\in\mathcal{A}}r_{\textsc{A}_{i}}(x;h)$ with random tie-breaking.
Implementations of Modiste
The human decision-making process with various forms of support can be effectively modeled as a stochastic contextual bandit problem. In this model, the diverse forms of support represent the available arms, and ${\mathcal{X}}$ represents the context space. For our investigation, we leverage two simple and efficient techniques from the existing contextual bandit literature: LinUCB (Li et al. 2010) and KNN-UCB (Guan and Jiang 2018). It’s worth noting that while any contextual bandit algorithm could be employed, we chose these two simpler methods as they have demonstrated their effectiveness in empirically showcasing the utility of our online decision support policy learning framework. We present specific implementations of the Modiste algorithm (see Algorithm 1) using LinUCB and KNN-UCB in Algorithms 2 and 3, respectively.
Algorithm 2 Modiste using LinUCB
1:  Input: human decision-maker $h$; UCB parameter $\alpha$
2:  Initialization: data buffer ${\mathcal{D}}_{0}=\{\}$; $\mathbf{A}_{a}=\mathbf{I}_{p\times p}$ and $\mathbf{b}_{a}=\mathbf{0}_{p\times 1}$ for all $a\in\mathcal{A}$; $\{\theta_{a}=(\mathbf{A}_{a})^{-1}\mathbf{b}_{a}:a\in{\mathcal{A}}\}$; human prediction error values $\{\widehat{r}_{a,0}(x;h)=\left\langle{\theta_{a}},{x}\right\rangle:x\in{\mathcal{X}},a\in{\mathcal{A}}\}$; initial policy $\pi_{1}(x)_{a}=1/\left|{\mathcal{A}}\right|$ for all $a\in{\mathcal{A}}$
3:  for $t=1,2,\dots,T$ do
4:     data point $(x_{t},y_{t})\in{\mathcal{X}}\times{\mathcal{Y}}$ is drawn iid from ${\mathcal{P}}$ (normalized s.t. $\left\lVert x_{t}\right\rVert_{2}\leq 1$)
5:     support $a_{t}\in{\mathcal{A}}$ is selected using policy $\pi_{t}$
6:     human makes the prediction $\widetilde{y}_{t}$ based on $x_{t}$ and $a_{t}$
7:     human incurs the loss $\ell(y_{t},\widetilde{y}_{t})$
8:     update the buffer ${\mathcal{D}}_{t}\leftarrow{\mathcal{D}}_{t-1}\cup\left\{{(x_{t},a_{t},\ell(y_{t},\widetilde{y}_{t}))}\right\}$
9:     update the decision support policy:
$\mathbf{A}_{a_{t}}~{}\leftarrow~{}$ $\mathbf{A}_{a_{t}}+x_{t}x_{t}^{\top}$
$\mathbf{b}_{a_{t}}~{}\leftarrow~{}$ $\mathbf{b}_{a_{t}}+r_{t}x_{t}$
$\theta_{a}~{}\leftarrow~{}$ $(\mathbf{A}_{a})^{-1}\mathbf{b}_{a}\text{ for all }a\in{\mathcal{A}}$
$\widehat{r}_{a,t}(x;h)~{}\leftarrow~{}$ $\left\langle{\theta_{a}},{x}\right\rangle\text{ for all }a\in{\mathcal{A}}$ (Step 1)
$\pi_{t+1}(x)~{}\leftarrow~{}$ $\operatorname*{arg\,min}_{a\in\mathcal{A}}\widehat{r}_{a,t}(x;h)-\alpha\cdot\sqrt{x^{\top}(\mathbf{A}_{a})^{-1}x}$ (Step 2)
10:  end for
11:  Output: policy $\pi^{\text{alg}}\leftarrow\pi_{T+1}$
Algorithm 3 Modiste using online KNN
1:  Input: human decision-maker $h$; warm-up steps $W$; number of neighbours $K$; exploration parameter $\gamma$
2:  Initialization: data buffer ${\mathcal{D}}_{0}=\{\}$; human prediction error values $\{\widehat{r}_{a,0}(x;h)=0.5:x\in{\mathcal{X}},a\in{\mathcal{A}}\}$; initial policy $\pi_{1}(x)_{a}=1/\left|{\mathcal{A}}\right|$ for all $a\in{\mathcal{A}}$
3:  for $t=1,2,\dots,T$ do
4:     data point $(x_{t},y_{t})\in{\mathcal{X}}\times{\mathcal{Y}}$ is drawn iid from ${\mathcal{P}}$
5:     support $a_{t}\in{\mathcal{A}}$ is selected using policy $\pi_{t}$
6:     human makes the prediction $\widetilde{y}_{t}$ based on $x_{t}$ and $a_{t}$
7:     human incurs the loss $\ell(y_{t},\widetilde{y}_{t})$
8:     update the buffer ${\mathcal{D}}_{t}\leftarrow{\mathcal{D}}_{t-1}\cup\left\{{(x_{t},a_{t},\ell(y_{t},\widetilde{y}_{t}))}\right\}$
9:     update the decision support policy:
${\mathcal{N}}(x)~{}\leftarrow~{}$ K neighbouring data points for $\text{K neighbouring data points for }x\text{ in }{\mathcal{D}}_{t}$
${\mathcal{N}}_{a}(x)~{}\leftarrow~{}$$\left\{{(x_{i},a_{i},\ell(y_{i},\widetilde{y}_{i})):(x_{i},a_{i},\ell(y_{i},\widetilde{y}_{i}))\in{\mathcal{N}}(x)\text{ and }a_{i}=a}\right\}\quad\text{ for all }a\in{\mathcal{A}}$
$\widehat{r}_{a,t}(x;h)~{}\leftarrow~{}$ $\frac{1}{\left|{\mathcal{N}}_{a}(x)\right|}\cdot\sum_{(x_{i},a_{i},\ell(y_{i},\widetilde{y}_{i}))\in{\mathcal{N}}_{a}(x)}{\ell(y_{i},\widetilde{y}_{i})}\quad\text{ for all }a\in{\mathcal{A}}\text{ with }\left|{\mathcal{N}}_{a}(x)\right|>0$ (Step 1)
[…]
$\pi_{\text{knn}}(x)~{}\leftarrow~{}$ $\operatorname*{arg\,min}_{a\in\mathcal{A}}\widehat{r}_{a,t}(x;h)$
$\pi_{t+1}(x)~{}\leftarrow~{}$ $\pi_{\text{rand}}(x)\quad\text{ if }t\leq W$ (Step 2)
$\pi_{t+1}(x)~{}\leftarrow~{}$ $\gamma\cdot\pi_{\text{rand}}(x)+(1-\gamma)\cdot\pi_{\text{knn}}(x)\quad\text{ if }t>W$ (Step 2)
10:  end for
11:  Output: policy $\pi^{\text{alg}}\leftarrow\pi_{T+1}$
Appendix C Additional Experimental Details
Figure 6: Examples of the Modiste interface for each form of support in CIFAR-$3A$ (Top) and MMLU-$2A$ (Bottom). From left to right, top row: Human Alone, Model, and Consensus; bottom row: Human Alone and LLM.
Datasets
Computational Only.
Two tasks (Synthetic-$2A$ and CIFAR-$2A$) are used only in the computational experiments, so we describe them abstractly. Both have the same set-up, but have different underlying data generating distributions. Consider learning a decision support policy in a setting where there are two forms of support available: $\mathcal{A}=\{\textsc{A}_{1},\textsc{A}_{2}\}$, hence CIFAR-$2A$. We simulate a decision-maker’s behavior under each $r_{\textsc{A}_{i}}$ across 3 classes for Synthetic/CIFAR. We instantiate the $m$ decision-makers in the population data using human simulations, where we randomly sample $r_{\textsc{A}_{i}}(x;h)$ from a distribution for each $h$ in the population.
Figure 7: We depict the latent space of the CIFAR subset used. Embeddings without filtering (left) and after filtering out points that violate separability of classes (right). We use the embeddings on the right for our Synthetic-$2A$ task and the left for the natural CIFAR tasks.
CIFAR.
To explore the performance of our algorithms in both separable and non-separable settings, we construct two subvariants of the CIFAR-10 (Krizhevsky 2009) image dataset. As depicted in Figure 7, we subsample from the CIFAR embeddings to create linearly separable classes. Embeddings are constructed by running t-SNE on the 512-dimensional latent codes extracted by the penultimate layer of a variant of the VGG architecture (VGG-11) (He et al. 2016) trained on the animal class subset of CIFAR; the model attained an accuracy of 89.5%. We filter the original 10 CIFAR classes to only include those involving animals (Birds, Cats, Deers, Dogs, Frogs, and Horses). We consider at most 5 such classes in our experiments (dropping Dog).
MMLU
We consider questions from four topics of MMLU (elementary mathematics, high school biology, high school computer science, and US foreign policy). Topic names match those proposed in (Hendrycks et al. 2020). We select topics to cover span an array of disciplines across the sciences and humanities, as discussed in Appendix E. We use questions in the MMLU test set for each topic that are at most 150 characters long. We limit the length of questions to facilitate readability in our human user studies and wanted to maximize parity between our computational and human experiments by considering the same set of questions across both set-ups. This yields 264, 197, 98, and 47 questions for the elementary mathematics, high school biology, US foreign policy, and high school computer science topics, respectively. We emphasize that in CIFAR, our forms of support operate in label space – in contrast, with MMLU, support selection is in covariate (topic) space – further highlighting the flexibility of our adaptive support paradigm.
We leverage OpenAI’s LLM-based embedding model (text-embedding-ada-003), to produce embeddings over the question prompts. Embedding vectors extracted via OpenAI’s API are by default length 1536; we apply t-SNE, like in the CIFAR tasks, to compress the embeddings into two-dimensional latent codes per example. Nicely, these latents are already separable (see Figure 5; no post-filtering is applied to ensure separability).
CIFAR Task Set-up
We consider a 5-class subset of CIFAR (Bird, Cat, Deer, Frog, and Horse), as discussed in Appendix D [^4] We explored a 3 class variant for HSEs to directly match the computational experiments; however, we realized that participants were able to figure out that which classes were impoverished, raising the base rate of correctly categorizing such images. Such behavior again highlights the need to carefully consider real human behavior in adaptive decision support systems.. We instantiate two forms of support: 1) a simulated AI model which provides a prediction for the image class, and 2) a consensus response derived from real humans, derived from the approximately 50 human annotators from CIFAR-10H (Peterson et al. 2019; Battleday, Peterson, and Griffiths 2020), presented as a distribution over the 5 classes.
We treat the original CIFAR-10 test set labels as the “true” labels and only include images for which the original CIFAR-10 label matches the label deemed most likely from the CIFAR-10H annotators (discrepancy was a rare occurrence, only 1.1% of the 5000 images considered). We sample a pool of 300 such images, and construct three different batches of 100 images. Participants are assigned to one batch of 100 images. Images per batch are sampled in a class-balanced fashion (i.e., 20 images for each of the 5 categories). Images are shuffled for each participant. The same latent codes ($z_{t}$) used in the computational experiments are employed for the user studies.
Corruptions and Support Design
We want to study whether our algorithms can properly learn which forms of decision support are actually needed by real humans. We therefore need to ensure that the task is sufficiently rich such that humans do need support (and the existing forms of support which can be of value). To mimic such a setting in CIFAR, we deliberately corrupt images of all classes except one (i.e., Birds [^5] We further disambiguate the non-corrupted bird class by upsampling to 160x160 images via Lanczos-upsampling following (Peterson et al. 2019; Battleday, Peterson, and Griffiths 2020; Collins, Bhatt, and Weller 2022) before presenting them to participants.) when presenting images to the user. Corruptions are formed via a composition of natural adversarial transformations proposed in Hendrycks et al. (2020), specifically shot noise ontop of glass blur. This enables us to check that our algorithms are able to properly recommend support for all other classes, as a human will be unable to decide on image category unassisted. Example corrupted images can be seen in Figure 3.
To ensure that the available forms of support have different regions of strength, we enforce that each form of support is only good at two of the five classes. In the case of the simulated AI model, we return the “true” class whenever the image is a Deer or Cat classes, and return one of the incorrect classes for all other images. For the consensus labels, we use the CIFAR-10H distribution derived from 50 annotators when the image is a Horse or Frog. As the CIFAR-10H labels were originally collected over all 10 CIFAR classes, sometimes an annotation was endorsed for a class outside of the 5 we consider; in that case, we discard the annotations assigned to those classes and renormalize the remaining distribution (on average, $<2\%$ of the original labeling mass was discarded per image). For all other classes, we intend the consensus to be an unhelpful form of support, as such, for images in the Bird, Deer, and Horse class, we return a uniform distribution (i.e., providing no information to the participant).
MMLU Task Set-up
Participants respond to 60 multiple choice questions, which were balanced by topic (15 questions per topic). Participants are assigned to one of three possible batches of 60 such questions. Order is shuffled. Participants are informed of the topic associated with each question (e.g., that the question was about biology). We implement a 10 second delay between when the question is presented and when the participant is allowed to submit their response to encourage participants to try each problem in earnest.
Topic Selection
We ran several pilot studies with Modiste to determine people’s base performance on a subset of MMLU topics. We intended to select a diverse set of topics such that it was unlikely any one participant would excel at all topics – as many of the topics were specialized and challenging (Hendrycks et al. 2020) – but also varied enough that participants may be strong in at least one area. We found that a large number of participants achieved reasonable performance on elementary mathematics, and that a sufficient number of participants also excelled at questions in the high school biology, US foreign policy, and high school computer science topics – though usually not all together.
We also factored in InstructGPT-3.5’s performance while deciding on the topics. To check whether our adaptive decision support algorithms are effective at learning good policies – like with CIFAR – it is helpful to have support available that is effective, should someone be unable to answer adequately alone. As a result, we looked more sympathetically on categories where model accuracy was already high.
However, the real-world is not so perfect: there may not be high-performing support available when a human struggles to make a decision. As such, we also deliberately forced down the accuracy of the LLM form of support in the mathematics topic. While we expect that most participants would be able to solve elementary mathematics problems without the aid of the LLM support, should they be unable to, the LLM was not able to help them. We leave further impovershing studies which mimic real-world support settings for future work.
Question Selection and Model Accuracy
The resulting accuracy of the LLM form of support on the questions shown to humans per topic is 29% for elementary mathematics, 89% for high school biology, 87% for US foreign policy, and 91% for high school computer science. We upweighted selection of foreign policy questions that participants in our pilot had gotten correct when constructing the question subset, as we wanted to ensure that, should a participant have foreign policy experience, they would be able to answer them.
Compute Resources
All computational experiments were run using CPUs (either on local machines or Google Colab), with the exception of training the VGG used for CIFAR embeddings. Here, we accessed a group compute cluster and trained the model for roughly five hours on a Nvidia A100-SXM-80GB GPU.
Appendix D Additional Computational Experiments
Figure 8: Snapshots of the learned decision support policies computed at the end of $100$ interactions with randomly sampled users for CIFAR-$3A$. The forms of support are colored in t-SNE embedding space. Both LinUCB and KNN generally identify the optimal form of support for a given input. All participant plots are included in Figure 12.
Table 2: Vary exploration parameters ($\alpha$ for LinUCB and $\gamma$ for KNN). We find that LinUCB generally performs better when exploration increases and KNN generally performs better when exploration decreases.
Algorithm
$r_{\textsc{A}_{1}}=[0.7,0.1,0.7]$
$r_{\textsc{A}_{2}}=[0.1,0.1,0.1]$
$r_{\textsc{A}_{1}}=[0.7,0.1,0.7]$
$r_{\textsc{A}_{2}}=[0.1,0.7,0.1]$
$r_{\textsc{A}_{1}}=[0.7,0.1,0.7]$
$r_{\textsc{A}_{2}}=[0.7,0.7,0.1]$
Synthetic-$2A$ Modiste-LinUCB $\alpha=0.1$ $0.14\pm 0.04$ $0.19\pm 0.09$ $0.34\pm 0.10$
Modiste-LinUCB $\alpha=10$ $0.13\pm 0.04$ $0.12\pm 0.03$ $0.30\pm 0.05$
Modiste-KNN $\gamma=0.01$ $0.12\pm 0.04$ $0.19\pm 0.08$ $0.33\pm 0.04$
Modiste-KNN $\gamma=0.2$ $0.16\pm 0.05$ $0.21\pm 0.06$ $0.34\pm 0.06$
CIFAR-$2A$ Modiste-LinUCB $\alpha=0.1$ $0.15\pm 0.08$ $0.14\pm 0.05$ $0.31\pm 0.05$
Modiste-LinUCB $\alpha=10$ $0.14\pm 0.04$ $0.18\pm 0.04$ $0.32\pm 0.05$
Modiste-KNN $\gamma=0.01$ $0.12\pm 0.03$ $0.14\pm 0.06$ $0.35\pm 0.06$
Modiste-KNN $\gamma=0.2$ $0.17\pm 0.05$ $0.20\pm 0.09$ $0.35\pm 0.06$
Algorithm
$r_{\textsc{A}_{1}}=[0.8,0.8,0.1,0.1]$
$r_{\textsc{A}_{2}}=[0.2,0.1,0.1,0.2]$
$r_{\textsc{A}_{1}}=[0.2,0.8,0.1,0.1]$
$r_{\textsc{A}_{2}}=[0.8,0.1,0.1,0.2]$
$r_{\textsc{A}_{1}}=[0.2,0.5,0.5,0.8]$
$r_{\textsc{A}_{2}}=[0.8,0.1,0.1,0.8]$
MMLU-$2A$ Modiste-LinUCB $\alpha=0.1$ $0.26\pm 0.06$ $0.22\pm 0.09$ $0.47\pm 0.08$
Modiste-LinUCB $\alpha=10$ $0.15\pm 0.05$ $0.18\pm 0.07$ $0.40\pm 0.06$
Modiste-KNN $\gamma=0.01$ $0.16\pm 0.05$ $0.15\pm 0.06$ $0.39\pm 0.06$
Modiste-KNN $\gamma=0.2$ $0.15\pm 0.03$ $0.12\pm 0.03$ $0.38\pm 0.05$
Varying KNN Parameters.
While LinUCB does not require identifying additional parameters, KNN has two: $K$ which is the number of nearest neighbors to select when estimating the risk of a form of support and $W$ which is the length of the warm-up period. In Table 3, we show that as long as $K$ is reasonably sized (i.e. $K>3$), the performance of KNN does not vary too much across datasets.
Table 3: We vary the $K$ and $W$ parameters used to instantiate KNN across three datasets and report the expected risk $L_{h}(\pi)$ (lower is better) across 5 runs.
$K=3$, $W=10$ $K=5$, $W=10$ $K=5$, $W=25$ $K=8$, $W=40$
Synthetic-$2A$ $0.20\pm 0.11$ $0.15\pm 0.05$ $0.14\pm 0.05$ 0.12 $\pm$ 0.04
CIFAR-$2A$ $0.25\pm 0.15$ $0.14\pm 0.04$ $0.17\pm 0.04$ $0.14\pm 0.04$
MMLU-$2A$ $0.18\pm 0.07$ $0.14\pm 0.04$ $0.15\pm 0.03$ $0.15\pm 0.04$
Table 4: On the CIFAR-$2A$ dataset, we explore the effect of varying the embedding size for KNN, with $K=8$, $W=25$ (Top), and LinUCB (Bottom). We report the expected risk $L_{h}(\pi)$ (lower is better) averaged over 5 runs. Recall that in our main paper experiments we used two-dimensional t-SNE embeddings, not these model embeddings, for ease of visualization.
Embedding
Size
$r_{\textsc{A}_{1}}=[0.7,0.1,0.7]$
$r_{\textsc{A}_{2}}=[0.1,0.7,0.1]$
$r_{\textsc{A}_{1}}=[0.7,0.1,0.7]$
$r_{\textsc{A}_{2}}=[0.7,0.7,0.1]$
$T=100$ $T=200$ $T=100$ $T=200$
2 $0.18\pm 0.08$ $0.14\pm 0.04$ $0.32\pm 0.05$ $0.33\pm 0.06$
4 $0.18\pm 0.04$ $0.16\pm 0.04$ $0.33\pm 0.05$ $0.33\pm 0.05$
8 $0.15\pm 0.05$ $0.15\pm 0.03$ $0.32\pm 0.05$ $0.33\pm 0.05$
512 $0.15\pm 0.05$ $0.15\pm 0.03$ $0.35\pm 0.05$ $0.35\pm 0.05$
Embedding
Size
$r_{\textsc{A}_{1}}=[0.7,0.1,0.7]$
$r_{\textsc{A}_{2}}=[0.1,0.7,0.1]$
$r_{\textsc{A}_{1}}=[0.7,0.1,0.7]$
$r_{\textsc{A}_{2}}=[0.7,0.7,0.1]$
$T=100$ $T=200$ $T=100$ $T=200$
2 $0.31\pm 0.05$ $0.32\pm 0.04$ $0.49\pm 0.05$ $0.49\pm 0.05$
4 $0.14\pm 0.04$ $0.14\pm 0.04$ $0.32\pm 0.04$ $0.31\pm 0.50$
8 $0.13\pm 0.04$ $0.13\pm 0.03$ $0.31\pm 0.03$ $0.31\pm 0.04$
512 $0.17\pm 0.03$ $0.16\pm 0.04$ $0.34\pm 0.04$ $0.33\pm 0.06$
Varying Embedding Size.
In the main text, we run computational experiments using a two-dimension t-SNE embedding. In Table 4 for CIFAR-$2A$, we consider how Modiste would behave when we vary dimensionality of $\mathcal{X}$. We extract higher dimensional embeddings from animal-class trained VGGs (see above) with varied penultimate layer widths, where width matches embedding size. While we opt for t-SNE embeddings, in the main paper, to provide a clearer visualizations, we find that strong performance holds even for high-dimensional embeddings. We expect that even larger embedding dimensions will permit Modiste to work in a variety of real-world contexts, but may come at the added cost of an increase in the number of required interactions, if we learn with no prior assumptions on the decision-maker’s expertise profile.
$r_{\textsc{A}_{1}}=[0.1,0.7,0.7]$
$r_{\textsc{A}_{2}}=[0.7,0.1,0.7]$
$r_{\textsc{A}_{3}}=[0.7,0.7,0.1]$
$r_{\textsc{A}_{1}}=[0.1,0.7,0.1]$
$r_{\textsc{A}_{2}}=[0.7,0.1,0.1]$
$r_{\textsc{A}_{3}}=[0.7,0.7,0.7]$
$r_{\textsc{A}_{1}}=[0.1,0.7,0.7]$
$r_{\textsc{A}_{2}}=[0.7,0.1,0.1]$
$r_{\textsc{A}_{3}}=[0.7,0.7,0.7]$
Figure 9: Loss plots over time for human simulations (denoted by $r_{\textsc{A}_{i}}$) in the CIFAR-$3A$ setting.
Table 5: We evaluate $|\mathcal{A}|>3$ via $L_{h}(\pi_{t})$ for $t=50,100,250$ for the same setup as Table 1, but with more arms. We find that Modiste works well over time. The success of LinUCB is highly dependent on the embedding space and its geometry.
Algorithm
$T=50$
$T=100$
$T=250$
CIFAR-$5A$ Modiste-LinUCB $0.59\pm 0.10$ $0.29\pm 0.17$ $0.34\pm 0.10$
Modiste-KNN $0.19\pm 0.15$ $0.14\pm 0.17$ $0.11\pm 0.15$
MMLU-$4A$ Modiste-LinUCB $0.13\pm 0.10$ $0.10\pm 0.07$ $0.17\pm 0.09$
Modiste-KNN $0.11\pm 0.12$ $0.08\pm 0.12$ $0.02\pm 0.04$
Figure 10: Snapshots of the learned decision support policies computed at the end of $60$ interactions with randomly sampled users for MMLU-$2A$ with LinUCB policies displayed atop KNN ones. The forms of support are colored in t-SNE embedding space. Topics correspond to clusters, clockwise from the top: math, biology, computer science, foreign policy.
$|\mathcal{A}|>3$ Experiments.
We now demonstrate how Modiste behaves when we increase the number of forms of support available to a decision-maker. Using the same set-up and algorithm-specific hyperparameters as Table 1, we convert the CIFAR setting to five forms of support (one for each class), resulting in CIFAR-$5A$, and convert the MMLU setting to four forms of support (one for each topic) resulting in MMLU-$4A$. In Table 5, we report $L_{h}(\pi)$ at three time steps: $t=50,100,250$. For CIFAR, we find that KNN significantly outperforms given more interactions, which is expected when we increase $|\mathcal{A}|$. We believe that LinUCB struggles as the t-SNE embedding space is likely not rich enough to capture the intricacies of all five forms of support. Given a larger embedding space, we expect that LinUCB would excel when the number of forms of support is increased. For MMLU, both LinUCB and KNN perform well by the 250th interaction: KNN is very good at lowering $L_{h}(\pi)$ on this task. In all, this experiment assures us that careful design of decision support can permit us to use Modiste when we have a larger set of potential forms of support.
While we include experiments with larger $|\mathcal{A}|$ here, we note that higher set sizes may also be problematic for two reasons: 1) decision-makers may struggle under many forms of support (Kalis, Kaiser, and Mojzisch 2013), and 2) as size increases, we find that the number of interactions required to learn an accurate personalized policy also increases.
Sensitivity to choice of $T$.
In Figure 9, we consider a setting of learning a policy with three forms of support: $\mathcal{A}=\{\textsc{A}_{1},\textsc{A}_{2},\textsc{A}_{3}\}$. This is an analogous setup to what we do in our human subject experiments in Section 7. We indicate the simulated human expertise profiles in the caption. We show how performance on a random sample of 100 points from a held-out set varies with time while learning a decision support policy. In general, we find that the more forms of support to learn, the longer it may take for the online algorithms to “converge.” For various simulations, KNN is more consistent in learning a decision support policy efficiently.
Appendix E Additional User Study Details and Results
Figure 11: Relationship between a participant’s sensibility of reliance (measured as the proportion of times they correctly agreed or disagreed with the form of support’s prediction) and the loss incurred by the participant’s learned policy. In the MMLU task, Modiste performs best for participants who sensibly rely on the provided support. Reliance is computed over all trials (100 for CIFAR-$3A$, 60 for MMLU-$2A$), and loss is averaged over the final 10 timesteps.
Figure 12: Snapshots of the recommended forms of support learned via Modiste for all participants in the CIFAR task. Policies learned using KNN get closer to optimal; see Figure 8.
Algorithm Parameters
For KNN, we take $K=8$, $W=25$, and $\epsilon=0.1$; for both algorithms, we let $\alpha=1$.  [^6] Due to a server-side glitch, 6 of the CIFAR 125 recruited participants received incorrect feedback on $\leq 2\%$ of trials. We use the same parameters for both tasks.
Further Results
Visualizing Learned Policies. We provide additional details on how the policy generalization snapshots in Figures 8 and 5 are constructed. We save out the parameters learned by each algorithm while a participant interacts with Modiste. After the 60 or 100 trials (depending on which task they participated in), we load in the final state of parameters. We sample the embeddings of unseen points from the respective task dataset (as described in Appendix C) and pass these through the respective algorithm (LinUCB or KNN, depending on which variant the participant was assigned to) – yielding a recommended form of support for said embedding. We color the latent by the form of support. We draw randomly 250 such unseen examples for CIFAR, and 100 for MMLU, when computing the policy recommendations.
We include snapshots for participants, across both tasks, in the main text. We depict all CIFAR examples and a random sampling of 7 of the 10 participants per variant for MMLU.
Participant Comments Hint at Mental Models of AI
In a post-experiment survey where we permitted participants to leave general comments, we observed that several participants in our MMLU task noted that their responses were biased by the AI. We include a sampling of comments provided by participants which we found particularly revealing. We believe that these responses motivate the need for further study of how users’ mental models of AI systems inform their decision-making performance, particularly in light of the rapidly advancing and public-facing foundation models (Bommasani et al. 2021).
  • ‘‘I thought the AI answers went against my previous notions about what an AI would be proficient at answering. I thought an AI would excel at anything where there is a set answer such as math, computer science coding question, and biology textbook definitions. Then I expected it to absoutly fail at political, and historical questions. Since they can be more opinionated and not have a definitive answer. However, the AI struggled with math and excelled in the other topics. So I was actively not choosing to agree with the AI on any math where I could also easily check its work and was heavily relying on it for everything else [...] I knew nothing about any question on computer science, so I mainly relied on AI for answers on that topic.’’
  • ‘‘AI seemed to me more helpful in topics I didn’t know, but sometime I trusted it too much, instead of thinking the answer through’’
  • ‘‘The AI had me second guessing myself sometimes so it affected my answers occasionally.’’
  • ‘‘Anything highlighted with the AI I was confident that would be the right answer.’’
Appendix F Further Discussion
Our human subject experiments validate trends observed in the computational experiments, demonstrating that Modiste can be used to learn decision support policies online. We find that learning personalized policies leads to performance benefits when encountering users with “varying” expertise profiles, and recovers fixed policies for participants with “strictly better” profiles. Future work modeling human idiosyncrasies (Steyvers and Kumar 2022) explicitly in synthetic decision-makers may bolster the effectiveness of personalization.
Participants’ reliance is one such factor. We define a metric for “reliance sensibility,” as the proportion of trials where the participant agreed with the support when correct and responded differently when the support was incorrect, out of all trials where the participant received support. The higher the proportion, the more “sensible” a participant’s degree of reliance on the provided support is, relative to what would be most beneficial for decision outcomes. [^7] Note that we cannot directly deduce whether, or to what degree, an individual participant relied on the form of support when presented - since we do not have access to what an individual would have said without support; alternative reliance inference schemes like (Tejeda et al. 2022) could be considered in future work. We plot the performance of the learned policy (i.e., incurred loss) against a participant’s reliance sensibility in Figure 11. In CIFAR-$3A$, participants largely had calibrated reliance because, by design, there was a signal for when to rely on each form of support. However, in the MMLU setting, we observe a variety of reliance behavior spanning over- and under-reliance, as we would expect in-the-wild (Zerilli, Bhatt, and Weller 2022). We also find a negative correlation between expected loss and reliance (Pearson $r$ correlation $=-0.47$ and $-0.59$ for KNN and LinUCB, respectively), indicating that Modiste was less effective for participants who did not appropriately rely on support. At present, Modiste only considers expertise explicitly when personalizing policies, yet these data underscore the importance of understanding when individual decision-makers may over-rely on potentially fallible support (Bussone, Stumpf, and O’Sullivan 2015). We explore practical considerations of personalized policies next.
An important concern of Modiste stems from our encouragement for individuals to disuse decision support. In particular, as machine learning models are increasingly used as a form of decision-support (Lai et al. 2023), our approach to learning a decision support policy promotes the disuse of models in favor of human judgment and agency. Automating access to model outputs may disadvantage individuals whose behavior (i.e., observed decisions) is misaligned with their values and may be used by organizations to prevent access to model outputs in a “big brother” sense (Zuboff 2023). We acknowledge that future work may need to justify to decision-makers when they are barred from accessing the model. This may lead to important intersections with the legal and policy communities. Instead of fully restricting access, less strict forms of disuse can be used to friction the decision-maker’s experience with a model, while having the same effect as barring access: examples include hiding the model output, which can be clicked on to appear (Buçinca, Malaya, and Gajos 2021), or displaying model confidence only for uncertain inputs (Antoran et al. 2020).
Additionally, since personalization inherently requires the use of participatory data (James et al. 2023), there will likely need to be explicit consent for the data to be used in learning decision support policies. Though our personalized policies make no assumptions about access to data from individuals a priori when learning a decision support policy, we use data collected from a like population, but we imagine that similar approaches can be used to design policies with offline user data (Charusaie et al. 2022; Buçinca et al. 2024). There may be unintended, negative impacts of collecting such data as organizations can use this data to implement nefarious decision support policies, whereby members are manipulated into using decision support, or to track user behavior. While we hope that personalized decision support policies encourage the thoughtful use of decision support, we caution against the ways that adverse actors can leverage these tools to privilege a select few, who in turn reap the benefits of personalized access to decision support tools (Gabriel et al. 2024).
Finally, the settings in which our tool, Modiste, was evaluated targeted English-language-speaking crowdworkers. The ecological validity (Zerilli, Bhatt, and Weller 2022) of our results may not hold for other groups, whose experiences, values, and preferences with respect to the use of machine learning models in decision-making may differ. While we did receive requisite approval to learn personalized policies on human subjects from our institution’s IRB, we hope that future work can cast a wider net for running experiments on groups of differing positionality (Kirk et al. 2024). Relatedly, our human subject experiments focused on two tasks (CIFAR and MMLU), which are relatively low-stakes decisions. We encourage future work to study similar problem formulations for personalized support in higher-stakes settings. We sincerely hope that decision support policies become a pervasive mechanism for empowering individuals to take control when models are provided to them.